OpenStudy (anonymous):

Write the expression cos(tan^-1)x) as an algebraic expression in x (without trig or inverse trig functions).

OpenStudy (anonymous):

There is a page about those in wikipedia and you can see it in the relationship table on the bottom: http://en.wikipedia.org/wiki/Inverse_trigonometric_functions However, if you look at this diagram http://en.wikipedia.org/wiki/File:Trigonometric_functions_and_inverse2.svg It's pretty easy to see that \[ tan(\theta) = \frac{x}{1} = x \] Now, the hypotenuse would be: \[ \text{hypotenuse} = \sqrt{1 + x^2} \] Now, if we do \[ tan^{-1}(x) = arctan(x) = \theta \] and now we can say: \[ cos(arctan(x)) = cos(\theta) = \frac{x}{hypotenuse} = \frac{x}{\sqrt{1+x^2}} \]